# Can LLMs Learn Macroeconomic Narratives from Social Media?

### v1

@@ Abstract
This study empirically tests the Narrative Economics hypothesis, which posits that narratives (ideas that are spread virally and affect public beliefs) can influence economic fluctuations. We introduce two curated datasets containing posts from X (formerly Twitter) which capture economy-related narratives. Employing Natural Language Processing (NLP) methods, we extract and summarize narratives from the tweets. We test their predictive power for macroeconomic forecasting by incorporating the tweets’ or the extracted narratives’ representations in downstream financial prediction tasks. Our work highlights the challenges in improving macroeconomic models with narrative data, paving the way for the research community to realistically address this important challenge. From a scientific perspective, our investigation offers valuable insights and NLP tools for narrative extraction and summarization using Large Language Models (LLMs), contributing to future research on the role of narratives in economics. [^1] Data will be shared upon paper acceptance.
@@ 1 Introduction
Existing evaluation strategies for narratives’ effectiveness for economic predictions are inadequate, focusing on macroeconomic correlations and anomalies, or on microeconomic predictions (§8). We propose utilizing downstream macroeconomic prediction tasks, where representations of latent narratives serve as input (§4-7). Interestingly, our results reveal that the successful narrative extraction we demonstrate offer only marginal improvement in macroeconomic prediction, compared to utilizing only financial information. In Section 9 we discuss the results’ implications on the validity of the Narrative Economics theory, at least when it comes to macroeconomics.
@@
  • Three curated tailored datasets: two Twitter collections and LLM-based analyses of tweets.
@@ 2.1 Twitter (X) Collected Datasets
Pre-Pandemic Twitter Dataset. We utilize Twitter API (Twitter, ) to collect a comprehensive dataset of 2.4 million tweets from January 2007 to December 2020, covering six topic areas with potential economic impact (economics, business, politics, current affairs, human disasters, and natural disasters) using targeted keywords detailed in App. A. For each topic, we iteratively retrieved tweets for each date within the timeframe. To prioritize viral tweets, we first retrieved the top 200 tweets based on follower count, then we randomly sampled 100 to mitigate potential bias towards highly active accounts typically associated with news outlets. This process yielded an average of roughly 400,000 tweets per topic, contributed by about 250,000 users per topic, each boasting an average follower count of 100 million, including global leaders, news outlets and other influencers. Minimal pre-processing was applied to the tweets, as common in literature (see §8 and App. A).
Post-2021 Twitter Dataset. This dataset, spanning September 2021 to July 2023, was specifically constructed for LLM-driven narrative analysis (see §4.2). To fairly test the predictive power of narratives, we needed to ensure that the employed LLM (Chat Completion API with GPT-3.5 (OpenAI, 2021), data cutoff September 2021) relies solely on provided tweets and pre-existing world knowledge, preventing access to ’future’ information. As our pre-pandemic collection concludes in December 2020, we assembled this new tweets dataset after the LLM’s knowledge cutoff date.
@@ 2.2 Macroeconomic Indicators
CBOE Volatility Index (VIX): Measures market expectations of future volatility based on S&P 500 options prices, often referred to as the ’fear gauge’ as it tends to rise during market stress and fall during market stability.
@@ 3 Experimental Setup
Synthetic ’narratives’: We generate synthetic narrative-like sentences expressing positive or negative cues, aligned with subsequent changes in the financial indicator. This allows us to assess the model’s ability to infer relationships between aligned narratives and the following market changes.
@@ 4.1 Financial Models
The financial models include traditional machine learning (ML) models (Linear Regression, Lasso, Ridge, SVM, Random Forest, and Logistic Regression), DA-RNN (Qin et al., 2017), and T5 (Raffel et al., 2020). Each model processes a sequence of historical financial values, either as individual features or as a time-series. T5 receives these inputs in textual form (’increase’ and ’decrease’) for ’direction change’ classification.
@@ 5.1 Tweets Narrative Analysis
Figure 2: Temporal distribution of top three narratives from the ’economics’ dataset, extracted by RELATIO (Ash et al., 2021) (see 8): UK’s Brexit, Greece’s financial debt, and Russia’s financial crises. We can see the evolving nature of these narratives over time, where the distribution is aligned with real-life related events.
@@ 5.2 LLM-Based Narrative Analysis
Having demonstrated the presence of narratives within our raw data, we now analyze the LLM’s ability to extract and summarize them. Following the methodology in § 4.2, we generated 697 LLM-based analyses, one for each date from September 2021 to July 2023. The outputs generated by the LLM can be divided into two main components: the analysis of tweets and the potential impact of them on the financial target. In each component, the analyses contain ~9 sentences on average.
@@ 6.1 Sentiment-Based Next-Day Prediction
Method. We use daily sentiment from the past week as features (see § 4.2) for FFR next-day prediction. Sentiments are fed as features into several classical ML models, described in § 2.2, applied to ’next value’ and ’direction change’ tasks. For the ’prediction change’ task the financial input is encoded either as numerical features or binary increase/decrease values to align with the label.
Results. Table 2 presents the results of the most performant models (full results are in App. G). In classifying ’direction change’, models with financial input (F & TF) outperform text-only models (T), exhibiting a 5% accuracy improvement (0.94 vs. 0.89). The slight difference between F and TF models (0.936, 0.939) suggests text has little impact. [^4] Feature importance confirms this, with the previous day’s financial feature dominating in TF models (not shown). The T models achieve comparable accuracy to the F baselines (0.89, 0.81). In predicting ’next value’, best F, TF and T models yield comparable MSE to the ’train-mean’ baseline (15.4, 15.6). [^5] Feature importance reveals low scores for both sentiments and financial features (not shown).
@@
Table 2: Results of most performant models for predicting FFR using sentiments. 1. In ’direction change’, financial features are encoded either numerically. TF and F perform nearly identically, outperforming T models and F baselines. 2. In ’next value’, best models are comparable to each other and to the F baseline.
@@ 6.2 Embeddings for Time-Series Prediction
Method. To better capture the richness and complexity of information concealed within the tweets, we turn to the two ’Embedding-Based Representations’ described in § 4.2. Additionaly, we step beyond ML models and instead utilize the time-series deep learning model DA-RNN (Qin et al., 2017), designed to capture temporal dynamics and complex relationships within data (see § 4.3).
Through rigorous evaluation, we explore various model configurations, target variables (FFR and VIX), tasks (’percentage change’, ’direction change’ and both together), prediction horizons, LLM architectures, aggregation methods, and the daily number of tweets. Lastly, we assess the models’ reliance on temporal context and relevant narratives, using the ’random texts’ and ’shuffled tweets’ counterfactual baselines described in § 3. Due to space constraints, we present results for a single task and setting- VIX ’next value’ prediction with 1- or 7- days horizons. Additional experiments showed a recurring pattern to the presented results.
Results. Figure 4 presents the models’ performance. For a 1-day horizon, the ’as-previous’ F baseline outperforms all other models (3.079 MSE). This suggests that the input information might not be beneficial for such a short-term prediction. For a 7-day horizon, both TF models (13.148, 13.147) initially appear to outperform the F model (13.463) and F baseline (16.172), implying a potential influence of the textual content. However, the ’random texts’ TF baseline outperforms (13.056) all other models, indicating that the good performance of the TF models is not likely due to the content of the tweets. We hypothesize that the presence of text improves performance, even when random, due to either spurious correlations or random noise that helps the model generalize, similarly to regularization techniques. The difficulty in capturing and representing the aggregated tweets information in a way that is meaningful for financial prediction might be a contributing factor. Additionally, challenges may lie in using historical financial data to predict future values of an indicator characterized by frequent random movements and fluctuations.
@@
Figure 4: VIX ’next value’ prediction for 1/7-days horizons. The F and TF baselines outperform all models in 1- and 7-day horizons, respectively, suggesting all models struggle to learn from tweets for the prediction.
@@ 7.1 LLMs for End-to-End prediction
Overall, the LLM consistently produced meaningful narrative analyses and comprehension of financial implications, although not being instructed to do so, but exhibited inconsistencies in integrating these insights for prediction. For example, it occasionally refused to provide predictions due to safety guardrails. In other cases it mirrored input ranges neglecting the potential impact of successfully analyzed narratives. When presented with ’synthetic narratives’ (§ 3), it recognized the change direction but struggled to quantify the magnitude of it. Addressing these limitations could unlock the full potential of LLMs for financial forecasting.
@@ 7.2 Two-Stage prediction with LLM Analyses
Method. The previous experiment revealed the LLM’s ability to create insightful analyses of tweets-financial data (see §5.2 and F.2 for LLM-outputs analysis). Here we use these analyses as inputs for a dedicated prediction model (see §4.2), to predict the S&P 500 ’direction change’. Unlike the embedding-based approach (§ 6.2) which struggled to aggregate diverse narratives, here the LLM produces concise analyses. Compared to the LLM predictor approach (§ 7.1), here we implement a separate fine-tuned model for the downstream prediction task. Additionally, we compare the models to the ’synthetic narratives’ T baseline (see § 3).
@@
Table 3: S&P 500 ’direction change’. TF and F models outperform all other models, with an insignificant gap of ~2% favoring the TF model.
@@ 10 Limitations
While our datasets were carefully curated to capture potential narratives, identifying them definitively is challenging, especially when aggregating multiple narratives for a holistic economic picture. The definition of ’narrative’ is broad and subjective, and narratives are typically only recognizable as such in retrospect. Combined with the inherent noise and susceptible to misinformation in social media, extracting reliable narratives with certainty is a complex task.
Lastly, we are limited to publicly accessible LLMs with a known cutoff date, to avoid potential ’future’ world knowledge. Utilizing other models might lead to better results.
@@ Appendix B Financial Indicators Details
The FFR is the interest rate at which depository institutions, such as banks, lend reserve balances to each other overnight to meet reserve requirements. The FFR serves as a critical monetary policy tool employed by the Federal Reserve to influence economic activity by adjusting the target range of the FFR in quarter-point increments following Federal Open Market Committee (FOMC) meetings. Changes in the FFR can have significant impacts on various sectors of the economy, including consumer spending, investments, housing, and financial markets. Changes in the FFR can have significant impacts on various sectors of the economy, including consumer spending, investments, housing, and financial markets, making it a widely monitored indicator of overall economic health. The dataset was downloaded from the publicly available economic data at FRED® website Fred . To address historical distribution changes, mainly due to the Great Recession and the COVID-19 pandemic, we employ ’blocking split for time-series data’, separating periods with consistent distributions. The longest block (December 2008-January 2015) yields 2207 samples and was used in experiments. Figure 5 shows the blocking splits of the FFR. The second block (December 2008-January 2015) is the longest and was used for experiments.
@@
The VIX measures market expectations of future volatility based on S&P 500 options prices. It reflects investors’ perceptions of the risk and uncertainty in the stock market over the next 30 days. The VIX is known as the ’fear gauge’ as it tends to rise during times of market stress and uncertainty and fall during periods of market stability. Analyzing changes in the VIX provides insights into investors’ sentiments, risk appetite, and economic confidence. This dataset was downloaded from Yahoo Finance® website (Yahoo Finance, ). Similar to the S&P 500, we utilize the close price of the index, either pre- or post-pandemic data, avoiding their integration within the same model.
@@ Appendix D LLM Details
Prompt: The prompt of the LLM contains instructions, a set of tweets from the analyzed month, and a set of the financial indicator values (VIX or S&P 500) from the same time period. We feed the LLM with both tweets and indicator values to allow it to potentially understand connections or correlations between them while generating the output. We limit to up to 10 tweets per day to avoid a long prompt. The sets of tweets and financial values are given in a dictionary format keyed by their date: ’yyyy-mm-dd: tweet or indicator value’. Each part of the prompt begins and ends with < > and </ > tags to enable easy post-processing. We instruct the model to generate a summarised analysis of the given input. We specify that the output will be used for predicting future indicator values. We explain about the input data and its structure. Finally, we break the task into two steps and instruct first to analyze the attributes that appear in the tweets and then analyze their potential effects on the close-future indicator values. A full example of a prompt is presented in Figure 6.

### v2

@@ Abstract
This study empirically tests the Narrative Economics hypothesis, which posits that narratives (ideas that are spread virally and affect public beliefs) can influence economic fluctuations. We introduce two curated datasets containing posts from X (formerly Twitter) which capture economy-related narratives. Employing Natural Language Processing (NLP) methods, we extract and summarize narratives from the tweets. We test their predictive power for macroeconomic forecasting by incorporating the tweets’ or the extracted narratives’ representations in downstream financial prediction tasks. Our work highlights the challenges in improving macroeconomic models with narrative data, paving the way for the research community to realistically address this important challenge. From a scientific perspective, our investigation offers valuable insights and NLP tools for narrative extraction and summarization using Large Language Models (LLMs), contributing to future research on the role of narratives in economics. [^1] Data is available through the Social Media Archive (SOMAR): https://socialmediaarchive.org/record/77.
@@ 1 Introduction
Existing evaluation strategies for narratives’ effectiveness for economic predictions are inadequate, focusing on macroeconomic correlations and anomalies, or on microeconomic predictions (§9). We propose utilizing downstream macroeconomic prediction tasks, where representations of latent narratives serve as input (§5-8). Interestingly, our results reveal that the successful narrative extraction we demonstrate offer only marginal improvement in macroeconomic prediction, compared to utilizing only financial information. In Section 10 we discuss the results’ implications on the validity of the Narrative Economics theory, at least when it comes to macroeconomics.
@@
  • Three tailored datasets: two Twitter curated collections and LLM-based analyses of tweets.
@@
2 Preliminaries
2.1 Macro- and Micro- Economics
Macroceconomics studies the behavior of the economy as a whole, examining factors like inflation, unemployment, and economic growth. Microeconomics, on the other hand, is concerned with the decision-making of individuals and firms, examining indicators like the price of a certain stock.
2.2 Definition of Narrative
The term "narrative" holds distinct meanings across disciplines. In Computational Narrative Understanding (CNU) (Chambers and Jurafsky, 2008, 2009), a narrative typically refers to a structured story with identifiable elements such as plot, characters, and setting, and the understanding of how these elements interrelate and contribute to the overall meaning and coherence of a narrative.
In contrast, our research operates within the definition of Narrative Economics. Within this framework, a narrative is a belief about the world that is shared by the population, regardless of its adherence to traditional narrative structures. This broader definition encompasses a wider range of communication phenomena, including tweets, opinions, and even rumors, that may lack the formal elements of a story in the CNU sense. These narratives can influence economic decisions and have a real affect in the world, even if they are not factually accurate. For instance, a narrative be the widespread belief that housing prices are increasing, whereas in reality (according to economic indicators) they are stagnating. This broader definition is particularly relevant to our study because we are interested in how widely held beliefs, structured in any way, can impact macroeconomic trends. Building on this foundation, such narratives can be identified and extracted using methods such as sentiment analysis or topic modeling, which analyze the aggregate sentiment and themes present in large collections of text data to reveal prevalent beliefs.
@@ 3.1 Twitter (X) Collected Datasets
Pre-Pandemic Twitter Dataset. We utilize Twitter API (Twitter, ) to collect a comprehensive dataset of 2.4 million tweets from January 2007 to December 2020, covering six topic areas with potential economic impact (economics, business, politics, current affairs, human disasters, and natural disasters) using targeted keywords detailed in App. A. For each topic, we iteratively retrieved tweets for each date within the timeframe. To prioritize viral tweets, we first retrieved the top 200 tweets based on follower count, then we randomly sampled 100 to mitigate potential bias towards highly active accounts typically associated with news outlets. This process yielded an average of roughly 400,000 tweets per topic, contributed by about 250,000 users per topic. The average follower count of these users was 100 million. This high average reflects our initial filtering strategy, which intentionally included users with very large followings, such as those of global leaders, news outlets, and other influential figures. Minimal pre-processing was applied to the tweets, as common in literature (see §9 and App. A).
Post-2021 Twitter Dataset. This dataset, spanning September 2021 to July 2023, was specifically constructed for LLM-driven narrative analysis (see §5.2). To fairly test the predictive power of narratives, we needed to ensure that the employed LLM (Chat Completion API with GPT-3.5 (OpenAI, 2021), data cutoff September 2021) relies solely on provided tweets and pre-existing world knowledge, preventing access to “future” information. As our pre-pandemic collection concludes in December 2020, we assembled this new tweets dataset after the LLM’s knowledge cutoff date.
@@ 3.2 Macroeconomic Indicators
CBOE Volatility Index (VIX): Measures market expectations of future volatility based on S&P 500 options prices, often referred to as the “fear gauge” as it tends to rise during market stress and fall during market stability.
@@ 4 Experimental Setup
Synthetic “narratives”: We generate synthetic narrative-like sentences expressing positive or negative cues, aligned with subsequent changes in the financial indicator. This allows us to assess the model’s ability to infer relationships between aligned narratives and the following market changes.
@@ 5.1 Financial Models
The financial models include traditional machine learning (ML) models (Linear Regression, Lasso, Ridge, SVM, Random Forest, and Logistic Regression), DA-RNN (Qin et al., 2017), and T5 (Raffel et al., 2020). Each model processes a sequence of historical financial values, either as individual features or as a time-series. T5 receives these inputs in textual form (“increase” and “decrease”) for “direction change” classification.
@@ 6.1 Tweets Narrative Analysis
Figure 2: Temporal distribution of top three narratives from the “economics” dataset, extracted by RELATIO (Ash et al., 2021) (see 9): UK’s Brexit, Greece’s financial debt, and Russia’s financial crises. We can see the evolving nature of these narratives over time, where the distribution is aligned with real-life related events.
@@ 6.2 LLM-Based Narrative Analysis
Having demonstrated the presence of narratives within our raw data, we now analyze the LLM’s ability to extract and summarize them. Following the methodology in § 5.2, we generated 697 LLM-based analyses, one for each date from September 2021 to July 2023. The outputs generated by the LLM can be divided into two main components: the analysis of tweets, and the potential impact of them on the financial target. In each component, the analyses contain ~9 sentences on average.
@@ 7.1 Sentiment-Based Next-Day Prediction
Method. We use daily sentiment from the past week as features (see § 5.2) for FFR next-day prediction. Sentiments are fed as features into several classical ML models, described in § 5.1, applied to “next value” and “direction change” tasks. For the “prediction change” task the financial input is encoded either as numerical features or binary increase/decrease values to align with the label.
Results. Table 2 presents the results of the most performant models (full results are in App. G). In classifying “direction change”, models with financial input (F & TF) outperform text-only models (T), exhibiting a 5% accuracy improvement (0.94 vs. 0.89). The slight difference between F and TF models (0.936, 0.939) suggests text has little impact. [^4] Feature importance confirms this, with the previous day’s financial feature dominating in TF models (not shown). The T models achieve comparable accuracy to the F baselines (0.89, 0.81). In predicting “next value”, best F, TF and T models yield comparable MSE to the “train-mean” baseline (15.4, 15.6). [^5] Feature importance reveals low scores for both sentiments and financial features (not shown).
@@
Table 2: Results of most performant models for predicting FFR using sentiments. 1. In “direction change”, financial features are encoded either numerically. TF and F perform nearly identically, outperforming T models and F baselines. 2. In “next value”, best models are comparable to each other and to the F baseline.
@@ 7.2 Embeddings for Time-Series Prediction
Method. To better capture the richness and complexity of information concealed within the tweets, we turn to the two “Embedding-Based Representations” described in § 5.2. Additionally, we step beyond ML models and instead utilize the time-series deep learning model DA-RNN (Qin et al., 2017), designed to capture temporal dynamics and complex relationships within data (see § 5.3).
Through rigorous evaluation, we explore various model configurations, target variables (FFR and VIX), tasks (“percentage change”, “direction change” and both together), prediction horizons, LLM architectures, aggregation methods, and the daily number of tweets. Lastly, we assess the models’ reliance on temporal context and relevant narratives, using the “random texts” and “shuffled tweets” counterfactual baselines described in § 4. These baselines are integrated with the F models, as explained in § 5.3, to create TF baselines.
Due to space constraints, we present results for a single task and setting- VIX “next value” prediction with 1- or 7- days horizons. Additional experiments showed a recurring pattern to the presented results.
Results. Figure 4 presents the models’ performance. For a 1-day horizon, the “as-previous” F baseline outperforms all other models (3.079 MSE). This suggests that the input information might not be beneficial for such a short-term prediction. For a 7-day horizon, both TF models (13.148, 13.147) initially appear to outperform the F model (13.463) and F baseline (16.172), implying a potential influence of the textual content. However, the “random texts” TF baseline outperforms (13.056) all other models, indicating that the good performance of the TF models is not likely due to the content of the tweets. We hypothesize that the presence of text improves performance, even when random, due to either spurious correlations or random noise that helps the model generalize, similarly to regularization techniques. The difficulty in capturing and representing the aggregated tweets information in a way that is meaningful for financial prediction might be a contributing factor. Additionally, challenges may lie in using historical financial data to predict future values of an indicator characterized by frequent random movements and fluctuations.
@@
Figure 4: VIX “next value” prediction for 1/7-days horizons. The F and TF baselines outperform all models in 1- and 7-day horizons, respectively, suggesting all models struggle to learn from tweets for the prediction.
@@ 8.1 LLMs for End-to-End prediction
Overall, the LLM consistently produced meaningful narrative analyses and comprehension of financial implications, although not being instructed to do so, but exhibited inconsistencies in integrating these insights for prediction. For example, it occasionally refused to provide predictions due to safety guardrails. In other cases it mirrored input ranges neglecting the potential impact of successfully analyzed narratives. When presented with “synthetic narratives” (§ 4), it recognized the change direction but struggled to quantify the magnitude of it. Addressing these limitations could unlock the full potential of LLMs for financial forecasting.
@@ 8.2 Two-Stage prediction with LLM Analyses
Method. The previous experiment revealed the LLM’s ability to create insightful analyses of tweets-financial data (see §6.2 and F.2 for LLM-outputs analysis). Here we use these analyses as inputs for a dedicated prediction model (see §5.2), to predict the S&P 500 “direction change”. Unlike the embedding-based approach (§ 7.2) which struggled to aggregate diverse narratives, here the LLM produces concise analyses. Compared to the LLM predictor approach (§ 8.1), here we implement a separate fine-tuned model for the downstream prediction task. Additionally, we compare the models to the “synthetic narratives” T baseline (see § 4).
@@
Table 3: S&P 500 “direction change”. TF and F models outperform all other models, with an insignificant gap of ~2% favoring the TF model.
@@ 11 Limitations
While our datasets were carefully curated to capture potential narratives, identifying them definitively is challenging, especially when aggregating multiple narratives for a holistic economic picture. The definition of “narrative” is broad and subjective, and narratives are typically only recognizable as such in retrospect. Combined with the inherent noise and susceptible to misinformation in social media, extracting reliable narratives with certainty is a complex task.
Regarding geographical aspects, this work is limited to English-language data, and to US-centric macroeconomic indicators.
Lastly, we are limited to publicly accessible LLMs with a known cutoff date, to avoid potential “future” world knowledge. Utilizing other models might lead to better results.
@@ Appendix B Financial Indicators Details
The FFR is the interest rate at which depository institutions, such as banks, lend reserve balances to each other overnight to meet reserve requirements. The FFR serves as a critical monetary policy tool employed by the Federal Reserve to influence economic activity by adjusting the target range of the FFR in quarter-point increments following Federal Open Market Committee (FOMC) meetings. Changes in the FFR can have significant impacts on various sectors of the economy, including consumer spending, investments, housing, and financial markets. Changes in the FFR can have significant impacts on various sectors of the economy, including consumer spending, investments, housing, and financial markets, making it a widely monitored indicator of overall economic health. The dataset was downloaded from the publicly available economic data at FRED® website Fred . To address historical distribution changes, mainly due to the Great Recession and the COVID-19 pandemic, we employ “blocking split for time-series data”, separating periods with consistent distributions. The longest block (December 2008-January 2015) yields 2207 samples and was used in experiments. Figure 5 shows the blocking splits of the FFR. The second block (December 2008-January 2015) is the longest and was used for experiments.
@@
The VIX measures market expectations of future volatility based on S&P 500 options prices. It reflects investors’ perceptions of the risk and uncertainty in the stock market over the next 30 days. The VIX is known as the “fear gauge” as it tends to rise during times of market stress and uncertainty and fall during periods of market stability. Analyzing changes in the VIX provides insights into investors’ sentiments, risk appetite, and economic confidence. This dataset was downloaded from Yahoo Finance® website (Yahoo Finance, ). Similar to the S&P 500, we utilize the close price of the index, either pre- or post-pandemic data, avoiding their integration within the same model.
@@ Appendix D LLM Details
Prompt: The prompt of the LLM contains instructions, a set of tweets from the analyzed month, and a set of the financial indicator values (VIX or S&P 500) from the same time period. We feed the LLM with both tweets and indicator values to allow it to potentially understand connections or correlations between them while generating the output. We limit to up to 10 tweets per day to avoid a long prompt. The sets of tweets and financial values are given in a dictionary format keyed by their date: “yyyy-mm-dd: tweet or indicator value”. Each part of the prompt begins and ends with < > and </ > tags to enable easy post-processing. We instruct the model to generate a summarized analysis of the given input. We specify that the output will be used for predicting future indicator values. We explain about the input data and its structure. Finally, we break the task into two steps and instruct first to analyze the attributes that appear in the tweets and then analyze their potential effects on the close-future indicator values. A full example of a prompt is presented in Figure 6.